voi at 3. A is completely dominant over a, B is completely dominant over b. Two alleles for a third trait, C and C are co-domPlease explain how to get the answer of 12, I got 27 and it is incorrect
Answer #1

AaBbC1C2 * AaBbC1C2

first let`s consider the cross

AaBb * AaBb

the number of possible phenotypes in dihybrid cross with the simple dominance relationship is 4,

number of phenotypes=2^ number of heterozygous loci= 2^2=4

for the co-dominant loci

C1C2 * C1C2

C1 C2
C1 C1C1 C1C2
C2 C1C2 C2C2

here C1C1, C1C2, C2C2 have different phenotypes, number of possible phenotypes=3

so total number of different phenotypes can be produced= number of phenotypes produced by AaBb * AaBb \times number of phenotypes produced by C1C2 * C1C2=4\times3=12

so the answer is c) 12